"The production function of a given firm Is given by Q =200sqrtLK and the.\nassociated cost function is given by 2L + 5K = 1000\n\nA. Calculate the level of labor and capital input at the optimum level of production\n\nB. Calculate the MRTSL, K and MRTS k, L\n\nc. What is the level of output at the least cost combination of labor and capital inputs?"


1
Expert's answer
2021-09-22T11:14:59-0400
"\frac{\delta Q}{\delta K}=100\sqrt{\frac{L}{K}}-5\lambda"

"\frac{\delta Q}{\delta L}=100\sqrt{\frac{K}{L}}-2\lambda"

"\lambda=20\sqrt{\frac{L}{K}}"

"L=2.5K"

"K=100"

"L=250"